# PromptPex: Automatic Test Generation for Language Model Prompts

### v1

@@ Abstract.
Large language models (LLMs) are being used in many applications and prompts for these models are integrated into software applications as code-like artifacts. These prompts behave much like traditional software in that they take inputs, generate outputs, and perform some specific function. However, prompts differ from traditional code in many ways and require new approaches to ensure that they are robust. For example, unlike traditional software the output of a prompt depends on the AI model that interprets it. Also, while natural language prompts are easy to modify, the impact of updates is harder to predict. New approaches to testing, debugging, and modifying prompts with respect to the model running them are required.
To address some of these issues, we developed PromptPex, an LLM-based tool to automatically generate and evaluate unit tests for a given prompt. PromptPex extracts input and output specifications from a prompt and uses them to generate diverse, targeted, and valid unit tests. These tests are instrumental in identifying regressions when a prompt is changed and also serve as a tool to understand how prompts are interpreted by different models. We use PromptPex to generate tests for eight benchmark prompts and evaluate the quality of the generated tests by seeing if they can cause each of four diverse models to produce invalid output. PromptPex consistently creates tests that result in more invalid model outputs than a carefully constructed baseline LLM-based test generator. Furthermore, by extracting concrete specifications from the input prompt, PromptPex allows prompt writers to clearly understand and test specific aspects of their prompts. The source code of PromptPex is available at https://github.com/microsoft/promptpex.
@@ 1. Introduction
Figure 1. Example illustrating that the use of PromptPex to test a given prompt against different models. For a given prompt (labeled System Prompt) and PromptPex-generated test input (on the left), the resulting output differs significantly depending on what AI model is used to interpret it. PromptPex automatically generates test input based on the prompt and evaluates whether the output is compliant with what the prompt specifies. Because the prompt specifies ”Return only the part of speech tag” the lower two models produced non-compliant output.
1.1. Motivation
@@
Large language models (LLMs) are being used in many applications beyond chatbots and prompts for these models are integrated into software applications as code-like artifacts. These prompts behave much like traditional software in that they take inputs, generate outputs, and perform some specific function (prompts-are-programs, ). They are also often part of complex chain of control flow that combines LLM-driven prompts and regular code supported by popular frameworks such as langchain (langchain, ).
Such prompts will become an integral part of many code bases in the future because they leverage the power of AI models and can perform common tasks such as summarization, classification, and evaluation that traditional non-AI software is unable to do. Furthermore, as AI models continue to diversify and become more efficient and effective, software that leverages them will benefit.
@@
While prompts are becoming a key element of software code bases, they have both similarities and differences from traditional software. Similarities include taking input, generating output, performing a transformation much like an ordinary function. But differences exist that create major new software engineering challenges. First, the output of a prompt is inherently non-deterministic due to the nature of the underlying AI model and inference engine that interprets it. Significant effort has been invested in ensuring the model output conforms, at least syntactically, to a given specification (openai-json, ). Also, while natural language prompts are easy for non-programmers to write, the effect of small changes to a prompt are unpredictable, leading to challenges making robust prompt edits.
@@
Second, unlike traditional software, where the behavior of a function depends only on the well-defined specification of the hardware that it is compiled to, the output of a prompt depends on the model that interprets it. As a result, if the AI model changes, the result of the prompt may change, sometimes dramatically (retain, ; prompt-migration-jahani2024generative, ). Application developers have strong motivations to change the underlying AI model their application uses because new models are being developed and released that are more efficient, more capable, able to run locally, available as open source, etc. As a result, it is critical that an application developer can quickly understand the implication of changing the underlying model on the behavior of a specific prompt. (Figure 1 illustrates a concrete example of this).
1.2. Our Solution: PromptPex
@@
While many new software engineering practices will need to be adapted to build robust AI software [^1] In this paper, we refer to AI software as any software that uses a generative AI model at runtime., in this paper we focus on helping developers test and evaluate individual prompts. Specifically our tool, PromptPex, takes a prompt as input and automatically generates and evaluates test cases that explore whether a given model interprets the prompt as directed. PromptPex was inspired by work on Program Exploration (Pex (tillmann2008pex, )), now shipping as IntelliTest in Microsoft Visual Studio (intellitest, ) [^2] Pex automatically generates unit tests for .Net applications using dynamic symbolic execution..
@@
From this extracted specification, we use an LLM to generate test cases focused on exploring whether the PUT with a given model (the Model Under Test, or MUT) adheres to the specification. We then run our test cases with different MUTs to generate model-specific outputs for a given test case, as shown in Figure 1. Because the tests were generated with the specification in mind, we can then automatically evaluate whether a given output from a MUT complies with the specification. While PromptPex cannot be used to automatically generate test cases that explore the full functionality of a given PUT, it is still valuable in creating test cases the break the requirements of our extracted specification. Details of how we extract the specifications and generate tests are provided in Section 3.
@@
Because it is fully automated, PromptPex is both easy to use and provides immediate insights into potential issues that can arise from the language in the PUT, the MUT, or the combination of the two. To evaluate our approach, we collected a suite of eight benchmark PUTs, created tests using both PromptPex and a baseline LLM-based test generator, ran each test with four MUTs, and evaluated the tests using an LLM to determine if the outputs were compliant with requirements specified in the PUT. We consider the generation of tests that are non-compliant as more successful because a test that causes a model to generate a non-compliant output indicates a problem that should be addressed. Our results show that PromptPex consistently generates more non-compliant tests than our baseline test generator and also clearly distinguishes relative model capabilities for the given prompt. A full explanation of our experiments and methods is available in Section 4.
1.3. Contributions
Our contributions are as follows:
@@ 1.3. Contributions
  • Systematic and Automated Test Generation for AI Model Prompts: While prior work has focused on optimizing a prompt for a given model (e.g.,  (prompt-opt-pryzant2023automatic, )) and generating unit test cases for programs  (molina2024test, ; xue2024llm4fin, ; zhang2024exploring, ; bhatia2024unit, ), our work is the first to focus on the specific problem of automated test generation for prompts and creating tests that allow a developer to understand the behavior of their prompt across multiple models.
@@
  • Specification Extraction from Prompts for Testing: To generate effective tests, we define a new approach to extract an input specification and output rules that capture targeted properties of prompts. These generated artifacts can be used both in automatically generating tests and in helping the prompt developer test, refine, and migrate their prompts to new models.
@@
  • Evaluation of Test Generation and Model Compliance with Specifications: We measure the effectiveness of PromptPex using a benchmark suite of eight benchmark prompts running the generated tests on four diverse AI models. We compare with a sophisticated baseline LLM-based test generator and show that our approach extracting a specification from the input prompt results in tests that are more likely to cause non-compliant outputs across all the models tested.
@@ 2. Motivating Example
The prompts we focus on, which are used as software artifacts, share some similarities but also have distinct features from chatbot prompts. First, prompts used in codebases act like programs, often with well-defined inputs and outputs. This may include both syntactic and semantic restrictions. While LLMs are generally good at handling inputs that do not adhere to format, the output must still be well-formed for other parts of the codebase, which may be traditional, to handle it effectively. These prompts exhibit constructs similar to various programming constructs, such as performing an early return for particular input, similar to an if-then-return in traditional programming. They can also include features like complex control flow, multiple returns, assertions, and constraints. All these features are described in natural language and can be easily modified or updated.
Unlike in chatbot applications, where users can often provide any input and the chatbot needs to handle it, the input domain of a prompt embedded inside a codebase is more narrowly defined. The prompts embedded in software expect well-defined input. The constraint over the input can be explicitly stated within the prompt itself to filter out invalid input or be implicit in the codebase so that invalid input is never sent to the prompt. These prompts do not operate in isolation. They are part of complex logic pipelines where the output from a prompt or a traditional program can be fed into one another. These pipelines can be thoroughly layered, made up of multiple prompts and programs processing data to create a single response. The input to a prompt may be preprocessed to optimize efficiency or to filter out invalid data. Likewise, the output can be validated, and if necessary, the prompt can be executed again with the same input for generating a well-formed response that can be correctly parsed by dependent components (TypeChat, ; pydanticValidatorsPydantic, ). The main advantage of using prompts embedded in software is their capability to perform tasks on natural language inputs, which is not possible with existing traditional programs.
The prompts that are part of codebases are rich in program-like constructs. In Appendix B.6, we describe one such prompt for extracting important entities from the text and returning them in a structured list format. Due to the complexity of that prompt, we use a simpler prompt as our running example as shown in Figure 2. Although this prompt is simple, it includes program-like constructs similar to those found in complex prompts and is representative in highlighting the challenges faced by the prompts being used in traditional software pipelines.
@@ 2.1. Prompt Under Test (PUT)
The prompt in Figure 2 is the part-of-speech classification (POS) prompt [^3] A modified version of the prompt from (prompt-opt-sammo-schnabel2024symbolic, )., which classifies a word into a part-of-speech tag. We have truncated the list of speech tags with their description for brevity. The complete prompt is available in Appendix B.1.
@@
  • Output: It defines the output as the part-of-speech tag, Unknown, or CantAnswer.
  • Computation: It describes how the output should be computed, which in this case is simply tagging the word as a part-of-speech tag.
@@
  • Assertions and constraints: Like in programs, various assertions and constraints can be defined directly in the prompt. However, they need not be explicitly implemented. For example, the tag must be from the list of tags provided.
@@
These prompts function like programs, featuring well-defined input and output specifications. For instance, when given the input quick brown fox jumps over the lazy dog; quick, the output generated is JJ. This input might originate from another prompt or program, and similarly, the output can be passed on to another component. While the part-of-speech prompt may align with what the prompt developer intended, ensuring that the LLM interpreting the prompt can precisely understand the intent and constraints remains a significant challenge. Developing these prompts is demanding because they require more rigorous testing than traditional software due to the vast potential input domain, even with constraints in place. Moreover, the model’s behavior can vary unpredictably based on the input, which can be neither fully understandable nor predictable. The inherent ambiguity of natural language further complicates efforts to manage these challenges effectively. LLMs often struggle to consistently follow provided instructions, but these instructions can be made more precise to accurately match the intended purpose, thereby minimizing unintended consequences. As models improve their ability to follow instructions, the alignment between the prompt developer’s intent and the model’s actual execution behavior will become increasingly critical.
@@
As an example of the challenges in writing an effective prompt, the POS prompt above sometimes generates a part-of-speech tag along with a description of the reasoning steps taken to arrive at that decision. This issue sometimes occurs even with gpt-4o, a state-of-the-art (SOTA) model. We observed this behavior due to ambiguity in what is allowed as output. To understand the ambiguity, note that the prompt requires a part-of-speech tag (e.g., NN) and not the word describing the part of speech (e.g., noun). Specifically, the prompt says ”Return only the part of speech tag.” The ambiguity arises because sometimes a model interprets this rule as applying only to tag (interpreting the prompt as ”Return only the part of speech tag and not the word describing the part of speech”) and does not explicitly forbid also adding an explanation about the reasoning behind the choice. This demonstrates the intricacies of prompt development and the value of both being more explicit about what is expected as well as testing model behavior more rigorously.
@@
The same prompt on gpt-4o most of the times only outputs the tag, but on gpt-3.5-turbo it often prefixes the tag with Output:, illustrating the problem of model portability. A prompt that works correctly on one model may not function as expected on another, highlighting the importance of extensive testing across multiple models. These challenges underscore the necessity for thorough testing of prompts on various models to identify issues early and to gain insights into how prompts behave during execution on different models. This understanding can help prompt developers address any problems and optimize their prompts for improved performance across different platforms.
@@
We developed PromptPex, a tool designed to help prompt developers better understand the execution behavior of their prompts. It achieves this by first extracting input and output specifications for the prompt under test (PUT), which are assertion-like constraints over the input and output, equivalent to pre- and post-conditions in a program.
PromptPex utilizes gpt-4o to extract these specifications from the prompt. The prompt developer can then examine these extracted specifications to compare their understanding with what a SOTA model perceives as expected input and the constraints on the output.
@@ 2.2. Input Specification (IS)
For the POS prompt, Figure 3 shows the extracted input specification (IS) by gpt-4o. For the PUT, the IS clearly captures that the input must be a sentence and a word, and it also lists constraints such as the requirement for the word to be a single word from the sentence. The single word constraint was not explicitly mentioned in the PUT, but during execution, the model must select a behavior, either allowing or disallowing compound words, such as ice cream. The IS highlights this under-specification in the PUT, indicating that the model has assumed it is expecting single words. Compound words might be considered valid input based on the developer’s understanding, but they would result in undefined behavior for the model and are more likely to lead to the output Unknown or CantAnswer.
@@
Figure 3. Extracted Input Specification for Part-of-Speech Prompt
@@ 2.3. Output Specification or Rules (OR)
The output specification, or the rules governing the output (OR), captures the constraints on the output generated for PUT. Figure 4 shows the extracted output specification (OR) for the POS prompt. The prompt developer can compare these rules with their understanding. For instance, the listed rules specify that the output must consist solely of the tag, without any additional text or formatting. This can be interpreted to mean that extraneous text, such as descriptions of the tag or formatting details, is not allowed—such as the use of Output: seen with gpt-3.5-turbo but it does not restrict description of the reasoning behind the tag. These rules help the prompt developer understand the potential behavior of the prompt during execution with the state-of-the-art model. This information can be used to tune the PUT to generate specifications that match more accurately the intentions of the prompt developer.
@@
Figure 4. Extracted Output Rules for Part-of-Speech Prompt

### v2

@@ Abstract.
Large language model prompts differ from traditional code in many ways and require new approaches to ensure that they are robust. For example, unlike traditional software the output of a prompt depends on the AI model that interprets it. Also, while natural language prompts are easy to modify, the impact of updates is harder to predict. New approaches to testing, debugging, and modifying prompts with respect to the model running them are required.
To address some of these issues, we developed PromptPex, an LLM-based tool to automatically generate and evaluate unit tests for a given prompt. PromptPex extracts input and output specifications from a prompt and uses them to generate diverse, targeted, and valid unit tests. These tests are instrumental in identifying regressions when a prompt is changed and also serve as a tool to understand how prompts are interpreted by different models. We use PromptPex to generate tests for twenty-two benchmark prompts and evaluate the quality of the generated tests by seeing if they can cause each of four diverse models to produce invalid output. PromptPex consistently creates tests that result in more invalid model outputs than a carefully constructed baseline LLM-based test generator. Furthermore, by extracting concrete specifications from the input prompt, PromptPex allows prompt writers to clearly understand and test specific aspects of their prompts. The source code of PromptPex is available at https://anonymous.4open.science/r/prompttest-83ED.
@@ 1. Introduction
Large language models (LLMs) are used in many applications beyond chatbots and prompts for these models are integrated into software applications as code-like artifacts. These prompts behave much like traditional software in that they take inputs, generate outputs, and perform a specific function (Guy et al., 2024). They are also often part of complex chain of control flow that combines LLM-driven prompts and regular code supported by popular frameworks such as langchain (16). Such prompts will become an integral part of many code bases in the future because they use the power of AI models and can perform common tasks such as summarization, classification, and evaluation that traditional software is unable to do. Furthermore, as AI models continue to diversify and become more efficient and effective, software that uses them will benefit.
Figure 1. Example illustrating the use of PromptPex to test a given prompt against different models. For a given prompt (labeled System Prompt) and PromptPex-generated test input (on the left), the resulting output differs depending on what AI model is used to interpret it. PromptPex automatically generates test input based on the prompt and evaluates whether the output is compliant with what the prompt specifies. Because the prompt specifies ”Return only the part of speech tag” the lower two models produced non-compliant output.
@@
While prompts are becoming a key element of software code bases, they have both similarities and differences from traditional software. Similarities include taking input, generating output, and performing a transformation much like an ordinary function. However, there are differences that create major new software engineering challenges. First, the output of a prompt is inherently non-deterministic due to the nature of the underlying AI model and inference engine that interprets it. Significant effort has been invested in ensuring the model output conforms, at least syntactically, to a given specification (OpenAI, 2024). Also, while prompts are easy for non-programmers to write, the effect of small changes to a prompt are unpredictable, leading to challenges making robust prompt edits.
@@
Second, unlike traditional software, where the behavior of a function depends only on the well-defined specification of the hardware that it is compiled to, the output of a prompt depends on the model that interprets it. As a result, if the AI model changes, the result of the prompt may change, sometimes dramatically (Dixit et al., 2024; Jahani et al., 2024). Application developers have strong motivations to change the underlying AI model their application uses because new models are being developed and released that are more efficient, more capable, able to run locally, available as open source, etc. As a result, it is critical that an application developer can quickly understand the implication of changing the underlying model on the behavior of a specific prompt. (Figure 1 illustrates a concrete example of this).
@@
Our Solution: PromptPex While many new software engineering practices will need to be adapted to build robust AI software [^1] We refer to AI software as any software that uses a generative AI model at runtime., in this paper we focus on helping developers test and evaluate individual prompts. Specifically our tool, PromptPex, takes a prompt as input and automatically generates and evaluates test cases that explore whether a given model interprets the prompt as directed.
@@
From this extracted specification, we use an LLM to generate test cases focused on exploring whether the PUT with a given model (the Model Under Test, or MUT) adheres to the specification. We then run our test cases with different MUTs to generate model-specific outputs for a given test case, as shown in Figure 1. Because the tests were generated with the specification in mind, we can then automatically evaluate whether a given output from a MUT complies with the specification. While PromptPex cannot be used to automatically generate test cases that explore the full functionality of a given PUT, it is still valuable in creating test cases that break the requirements of our extracted specification.
@@
PromptPex is both easy to use and provides immediate insights into potential issues that can arise from the language in the PUT, the MUT, or the combination of the two. To evaluate our approach, we collected a suite of twenty-two benchmark PUTs, created tests using both PromptPex and a baseline LLM-based test generator, ran each test with four MUTs, and evaluated the tests using an LLM to determine if the outputs were compliant with requirements specified in the PUT. We consider the generation of tests that are non-compliant as more successful because a test that causes a model to generate a non-compliant output indicates a problem that should be addressed. Our results show that PromptPex consistently generates more non-compliant tests than our baseline test generator and also clearly distinguishes relative model capabilities for the given prompt.
Contributions This paper makes the following contributions.
  1. (1)
@@
    Systematic and Automated Test Generation for AI Model Prompts: Our work is the first to focus on the specific problem of automated test generation for prompts and creating tests that allow a developer to understand the behavior of their prompt across multiple models. Prior work has focused on prompt optimization for a model (e.g.,  (Pryzant et al., 2023)) and generating unit test cases for programs  (molina2024test; xue2024llm4fin; zhang2024exploring; bhatia2024unit).
  2. (2)
@@
    Specification Extraction from Prompts for Testing: To generate effective tests, we define a new approach to extract an input specification and output rules that capture targeted properties of prompts. These generated artifacts can be used both to automatically generate tests and help the prompt developer test, refine, and migrate their prompts to new models.
  3. (3)
@@
    Evaluation of Test Generation and Model Compliance with Specifications: We measure the effectiveness of PromptPex using a benchmark suite of twenty-two benchmark prompts running the generated tests on four diverse AI models. We compare with a sophisticated baseline LLM-based test generator and show that our approach extracting a specification from the input prompt results in tests that are more likely to cause non-compliant outputs across all the models tested.
@@ 2. Motivating Example
The prompts we focus on, which are used as software artifacts, share some similarities but also have distinct features from chatbot prompts. First, prompts used in codebases act like programs, often with well-defined inputs and outputs. This may include both syntactic and semantic restrictions. While LLMs are generally good at handling inputs that do not adhere to format, the output must still be well-formed for other parts of the codebase, which may be traditional, to handle it effectively. These prompts exhibit constructs similar to programming constructs, such as performing an early return for particular input, similar to an if-then-return in traditional programming. They can also include features like complex control flow, multiple returns, assertions, and constraints. All these features are described in natural language and can be easily modified.
Unlike in chatbot applications, where users can often provide any input and the chatbot needs to handle it, the input domain of a prompt embedded inside a codebase is more narrowly defined. The prompts embedded in software expect well-defined input. The constraint over the input can be explicitly stated within the prompt itself to filter out invalid input or be implicit in the codebase so that invalid input is never sent to the prompt. These prompts do not operate in isolation. They are part of complex logic pipelines where the output from a prompt or a traditional program can be fed into one another. These pipelines can be thoroughly layered, made up of multiple prompts and programs processing data to create a single response. The input to a prompt may be preprocessed to optimize efficiency or to filter out invalid data. Likewise, the output can be validated, and if necessary, the prompt can be executed again with the same input to generate a well-formed response that can be correctly parsed by dependent components (37; 38).
The prompts that are part of codebases are rich in program-like constructs. We used a collection of such prompts in our evaluation. We use a simpler prompt as our running example as shown in Figure 2. Although this prompt is simple, it includes program-like constructs similar to those found in complex prompts and is representative in highlighting the challenges faced by the prompts being used in traditional software pipelines.
@@ 2.1. Prompt Under Test (PUT)
The prompt in Figure 2 is the part-of-speech classification (POS) prompt [^2] A modified version of the prompt from (Schnabel and Neville, 2024)., which classifies a word into a POS tag. We have truncated the list of speech tags with their description for brevity.
@@
  • Output: It defines output as POS tag, Unknown, or CantAnswer.
  • Computation: It describes how the output should be computed, which in this case is simply tagging the word as a POS tag.
@@
  • Assertions and constraints: Like in programs, assertions and constraints can be defined directly in the prompt. However, they need not be explicitly implemented. For example, the tag must be from the list of tags provided.
@@
These prompts function like programs, featuring well-defined input and output specifications. For instance, when given the input quick brown fox jumps over the lazy dog; quick, the output generated is JJ. This input might originate from another prompt or program, and similarly, the output can be passed on to another component. While the POS prompt may align with what the prompt developer intended, ensuring that the LLM interpreting the prompt can precisely understand the intent remains a challenge. Developing these prompts is demanding because they require more rigorous testing than traditional software due to the vast potential input domain, even with constraints in place. Moreover, the model’s behavior can vary unpredictably based on the input, which can be neither fully understandable nor predictable. LLMs often struggle to consistently follow provided instructions, but these instructions can be made more precise to accurately match the intended purpose, thereby minimizing unintended consequences. As models improve their ability to follow instructions, the alignment between the prompt developer’s intent and the model’s actual execution behavior will become increasingly critical.
@@
As an example of the challenges in writing an effective prompt, the POS prompt above sometimes generates a POS tag along with a description of the reasoning steps taken to arrive at that decision. This issue sometimes occurs even with gpt-oss, our most effective model-under-test. We observed this behavior due to ambiguity in what is allowed as output. To understand the ambiguity, note that the prompt requires a POS tag (e.g., NN) and not the word describing the part of speech (e.g., noun). Specifically, the prompt says ”Return only the part of speech tag.” The ambiguity arises because sometimes a model interprets this rule as applying only to tag (interpreting the prompt as ”Return only the part of speech tag and not the word describing the part of speech”) and does not forbid adding an explanation about the reasoning behind the choice. This demonstrates the intricacies of prompt development and the value of both being more explicit about what is expected and testing model behavior rigorously.
@@
The same prompt on gpt-oss most of the times only outputs the tag, but on gpt-3.5 it often prefixes the tag with Output:, illustrating the problem of model portability. A prompt that works correctly on one model may not function as expected on another, highlighting the importance of extensive testing across multiple models. These challenges underscore the necessity for thorough testing of prompts on various models to gain insights into how prompts behave during execution on different models. This understanding can help prompt developers address problems and optimize their prompts across different platforms.
@@
We developed PromptPex, a tool designed to help prompt developers better understand the execution behavior of their prompts. It achieves this by first extracting input and output specifications for the prompt under test (PUT), which are assertion-like constraints over the input and output, equivalent to pre- and post-conditions in a program. PromptPex utilizes gpt-oss to extract these specifications from the prompt. The prompt developer can examine these extracted specifications to compare their understanding with what a SOTA model perceives as expected input and output constraints.
@@ 2.2. Input Specification (IS)
For the POS prompt, Figure 3 shows the extracted input specification (IS) by gpt-oss. For the PUT, IS specifies that the input must be a sentence and a word, with constraints like the word being a single word from the sentence. Though not explicitly stated in PUT, the model must decide whether to allow compound words (e.g., ice cream). The IS highlights this under-specification in the PUT, indicating that the model has assumed it is expecting single words. Compound words might be considered valid input based on the developer, but they would result in undefined behavior for the model and are likely to lead to the output Unknown or CantAnswer.
@@
Figure 3. Extracted Input Specification for POS Prompt
@@ 2.3. Output Specification or Rules (OR)
The output specification, or the rules governing the output (OR), captures the output constraints generated for PUT. Figure 4 shows the extracted (OR) for the POS prompt. The prompt developer can compare these rules with their understanding. For instance, the listed rules specify that the output must consist solely of the tag, without any additional text or formatting. This can be interpreted to mean that extraneous text, such as descriptions of the tag or formatting details, is not allowed—such as the use of Output: seen with gpt-3.5-turbo but it does not restrict description of the reasoning behind the tag. These rules help the prompt developer understand the potential behavior of the prompt during execution and can be used to tune the PUT to generate specifications that more accurately match the intentions of the prompt developer.
@@
Figure 4. Extracted Output Rules for POS Prompt